From a group of 100 students, 30 are taking math, 20 are taking chemistry, and of those totals 10 are taking both. What is the probability that a randomly selected student is taking math or chemistry?

Answer by CharlesG2(834)

From a group of 100 students, 30 are taking math, 20 are taking chemistry, and of those totals 10 are taking both. What is the probability that a randomly selected student is taking math or chemistry?

let set A = all students = 100 students
let set M = math students = 30 students
let set C = chemistry students = 20 students
probability of just math = 30/100 = 30%
probability of just chemistry = 20/100 = 20%
probability of both math and chemistry = 10/100 = 10%
probability of math or chemistry but not both = (30 - 10)/100 + (20 - 10)/100
probability of math or chemistry but not both = 20/100 + 10/100 = 30/100 = 30%